In the study from the previous question, x=59 adults have been confirmed hypertensive out of n=200 participants. Assuming the standard normal distribution can be applied and p=0.313 under the null hypothesis, what is the Z-statistic for x=59? Recall that the expected value = np and the standard deviation = square root of n p open parentheses 1 minus p close parentheses end root .

= -0.55

Continuing from the previous question, if we observed 78 hypertensive adults in the 2nd random sample of 200, what is the one-sided p-value associated with x=78, given the prevalence of 0.313 under the null hypothesis? Here, we use the one-sided p-value = Pr(X > 78).

a. 0.024

b. 0.009

c. 1.98

d. -0.009
Answer #1

From the given information,

Mean=n*p=200*0.313=62.6

SD=6.5579

Hence,

P(X>78)=2.3483

Hence,

Required p-value is,

p-value= 0.009

i.e., option b. is correct.

Dear student,
I am waiting for your feedback. I have given my 100% to solve your queries. If you are satisfied by my given answer. Can you please like it☺
  
Thank You!!!